OpenStudy (anonymous):

WILL MEDAL A number is multiplied by 3. If 3 is added to the product, you get 6. What is the number?

OpenStudy (alexandervonhumboldt2):

a*3 a*3+3=6

OpenStudy (alexandervonhumboldt2):

first subtract 3 from both sides: 3a+3-3=6-3 3a=3 divide both sides by 3 3a/3=3/3 a=3/3=?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Thank you! (: